# 22footnotetext: To whom correspondence should be addressed

### v1

@@ Abstract
Irreducible Cartesian tensors (ICTs) play a crucial role in the design of equivariant graph neural networks, as well as in theoretical chemistry and chemical physics. Meanwhile, the design space of available linear operations on tensors that preserve symmetry presents a significant challenge. The ICT decomposition and a basis of this equivariant space are difficult to obtain for high-order tensors. After decades of research, we recently achieve an explicit ICT decomposition for $n=5$ (bonvicini2024irreducible) with factorial time/space complexity. This work, for the first time, obtains decomposition matrices for ICTs up to rank $n=9$ with reduced and affordable complexity, by constructing what we call path matrices. The path matrices are obtained via performing chain-like contraction with Clebsch-Gordan matrices following the parentage scheme. We prove and leverage that the concatenation of path matrices is an orthonormal change-of-basis matrix between the Cartesian tensor product space and the spherical direct sum spaces. Furthermore, we identify a complete orthogonal basis for the equivariant space, rather than a spanning set (pearce2023brauer), through this path matrices technique. We further extend our result to the arbitrary tensor product and direct sum spaces, enabling free design between different spaces while keeping symmetry. The Python code is available in the appendix where the $n=6,\dots,9$ ICT decomposition matrices are obtained in <0.1s, 0.5s, 1s, 3s, 11s, and 4m32s, respectively.
@@ 1 Introduction
The deployment of equivariant neural networks in physics, chemistry, and robotics often requires certain equivariance or invariance to effectively incorporate inductive bias. Equivariant Graph Neural Networks (EGNNs) are well-suited for this purpose, as they can maintain equivariance while processing features within the graph (zaverkin2024higher; simeon2024tensornet; batzner20223). One of the most promising approach to constructing equivariant layers involves leveraging irreducible representations. Depending on the basis of the space, two distinct types of irreducible representations are utilized in different EGNNs: spherical (spherical tensors) (musaelian2023learning; batatia2022mace) and Cartesian irreducible representations (Irreducible Cartesian Tensors, ICTs) (zaverkin2024higher; schutt2021equivariant). Spherical tensors serve as the starting point for EGNN based on irreducible representations. However, they require Clebsch-Gordan transforms with significant computational overhead to derive desired irreducible representations from existing ones during each inference, making them computationally intensive. In contrast, ICTs are more efficient, as high-order representations can be easily computed using tensor products (specifically, outer products in EGNNs), making them more computationally friendly for most deep learning frameworks. The ICTs are obtained through the ICT decomposition matrices $H$,
$T^{(n;j;q)}=H^{(n;j;q)}T^{(n)},$ (1)
where $T^{(n;j;q)}$ is an ICT, $T^{(n)}$ is a given tensor, $n$ represents the rank of the tensor, $j$ is called the weight of the irreducible representation, and $q$ denotes the label (since the multiplicity for given $n$ and $j$ is sometimes greater than $1$). The decomposition matrices $H$ should be constructed such that $T^{(n;j;q)}$ are irreducible. The challenge lies in the fact that $H$ of rank $n>5$ are unknown. The recent development of EGNNs research benefits from the irreducible decomposition matrices $H$ obtained from the physics and chemistry communities. The decomposition matrices for $n=4$ have been established by andrews1982irreducible. However, only very recently have we obtained the decomposition matrices for $n=5$ (bonvicini2024irreducible). Due to these limitations, state-of-the-art EGNNs cannot freely manipulate high-order ICTs or can only manipulate those whose weight equals their rank (zaverkin2024higher).
@@
Method Ranks Orthogonality Time/Space complexity
@@
Table 1: Comparison with previous works.
@@
This work has three main contributions:
  • We propose a method to efficiently calculate high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$.
@@
The following contents are organized as follows. In Section 2, we recall some basic ideas of representation theory, spherical harmonics, tensors, tensor product spaces, and we also introduce the Clebsch-Gordan coefficients and the parentage scheme. In Section LABEL:sec:problem, we formulate our problems: obtaining high-order ICT decomposition matrices and bases for arbitrary tensor product spaces. In Section LABEL:sec:highorderict, we introduce our theory and method to obtain high-order ICT decomposition matrices. In Section LABEL:sec:basis, we present the theory and method to obtain bases of equivariant spaces where the input and output spaces are the same. In Section LABEL:sec:extension, we extend the theory to the equivariant design spaces where the input and output spaces are different, and show the method to obtain bases for such general spaces. In Section LABEL:sec:relatedworks, we review related works on designing EGNNs and calculating ICT decomposition matrices. In Section LABEL:sec:discussion, we discuss the relation between pearce2023brauer and this work, as well as the theory for $SO(n)$, $O(n)$, and $SU(n)$. We provide concluding remarks in Section LABEL:sec:conclusion, and discuss the potential social impact in Section LABEL:sec:boardersocialimpact. We introduce the reproducibility details in Section LABEL:sec:repro. The acknowledgments are presented after the main text. The code and visualizations are shown in the appendix.
@@ 2 Preliminaries
2.1 Representations and group equivariance
We first recall some basic facts. A thorough introduction to group theory can be found in inui2012group and brocker2013representations. We first focus on the $O(3)$ group (3D rotation and reflection), but we also extend our conclusions to $O(n)$, $SO(n)$, and $SU(n)$ in Section LABEL:sec:discussion. Consider a vector $v\in V$. The rotation and reflection operations should be able to act on it. Representation theory shows that $O(3)$ can be represented by some real matrices in the Cartesian space via a group homomorphism $\rho:G\to\mathbbm{R}^{3\times 3}$, which is called the natural representation. It holds that
$\rho(g_{1}*g_{2})=\rho(g_{1})\cdot\rho(g_{2})$
for any $g_{1}$, $g_{2}\in G$. The rotation $g$ acting on the vector $v$ can therefore be represented as a matrix multiplication $\rho(g)\cdot v$. Sometimes, one may also refer to the space of $v$ as a representation space. If the results of the group action remain in the same space, i.e., $\rho(g)v\in V$. In this sense, this place is also called a $G$-invariant space or subspace. To represent $O(3)$, the matrix $\rho(g)$ is constrained to be orthogonal. It is important to note that the co-domain of $\rho$ can differ, meaning it can also be represented by matrices of other dimensions. In this paper, we denote the co-domain of $\rho$ using a subscript when it should be made explicit, e.g., $\rho_{\mathbbm{R}^{3}}$. Now, it suffices to define equivariance. We claim that a function $f:X\to Y$ is equivariant if and only if
$\rho_{Y}f(x)=f(\rho_{X}(x)).$
We only consider $f$ as a linear function, and thus it is represented as a matrix. The set of all such $f$ constitutes a vector space. We sometimes denote this space as ${\rm End}_{O(3)}(X)$ if $X=Y$, and as ${\rm Hom}_{O(3)}(X,Y)$ otherwise. Meanwhile, we say that the function $f$ is invariant if $\rho_{Y}=1$. Since $X$ is not necessarily the same as $Y$, this actually makes finding the function $f$ relatively straightforward. Specifically, to fulfill
$\rho_{Y}(g)M=M\rho_{X}(g),$
we can simply let
$\rho_{Y}(g)=M\rho_{X}(g)M^{-1}.$
However, this approach is very time-consuming in the design of EGNNs. The matrix $M$ here is learnable in the networks. During training, we must repeatedly compute the inverse of $M$, which is computationally expensive. Even if we could afford this, the results of the $M_{i}v_{i}$ is not addable in general, as the co-domains of different $M_{i}$ vary. Moreover, we cannot guarantee that $M$ is actually invertible without adding restriction on the determinants.
Considering that linear combinations involve a tremendous number of inversion operations, this process becomes intractable. Therefore, we first focus on the simpler case where $X=Y$. We will later generalize this to the case where $X\neq Y$, but $Y$ is known and remains consistent across different $M$. When it is necessary to emphasize the change in space, we refer to a linear mapping $M$ as $G$-$(X,Y)$-equivariant.
2.2 Irreducible representations and spherical spaces
We then introduce the definition of reducible and irreducible representation, in a matrix view:
Definition 1
A group homomorphism $\rho\in{\rm Hom}_{G}(\mathbbm{R}^{k})$ is said to be reducible, if there is an invertible matrix $P$, such that for every $g\in G$, we have
$P\rho(g)P^{-1}={\rm diag}\big{(}\rho_{1}(g),\rho_{2}(g)\big{)}$
for some non-trivial $\rho_{1}$ and $\rho_{2}$.
The $3\times 3$ rotation and reflection matrices are irreducible. More generally,

### v2

@@ Abstract
Irreducible Cartesian tensors (ICTs) play a crucial role in the design of equivariant graph neural networks, as well as in theoretical chemistry and chemical physics. Meanwhile, the design space of available linear operations on tensors that preserve symmetry presents a significant challenge. The ICT decomposition and a basis of this equivariant space are difficult to obtain for high-order tensors. After decades of research, we recently achieve an explicit ICT decomposition for $n=5$ (bonvicini2024irreducible) with factorial time/space complexity. This work, for the first time, obtains decomposition matrices for ICTs up to rank $n=9$ with reduced and affordable complexity, by constructing what we call path matrices. The path matrices are obtained via performing chain-like contraction with Clebsch-Gordan matrices following the parentage scheme. We prove and leverage that the concatenation of path matrices is an orthonormal change-of-basis matrix between the Cartesian tensor product space and the spherical direct sum spaces. Furthermore, we identify a complete orthogonal basis for the equivariant space, rather than a spanning set (pearce2023brauer), through this path matrices technique. We further extend our result to the arbitrary tensor product and direct sum spaces, enabling free design between different spaces while keeping symmetry. The Python code is available in https://github.com/ShihaoShao-GH/ICT-decomposition-and-equivariant-bases where the $n=6,\dots,9$ ICT decomposition matrices are obtained in 1s, 3s, 11s, and 4m32s, respectively.
@@ 1 Introduction
The deployment of equivariant neural networks in physics, chemistry, and robotics often requires certain equivariance or invariance to effectively incorporate inductive bias. Equivariant Graph Neural Networks (EGNNs) are well-suited for this purpose, as they can maintain equivariance while processing features within the graph (zaverkin2024higher; simeon2024tensornet; batzner20223). One of the most promising approach to constructing equivariant layers involves leveraging irreducible representations. Depending on the basis of the space, two distinct types of irreducible representations are utilized in different EGNNs: spherical (spherical tensors) (musaelian2023learning; batatia2022mace) and Cartesian irreducible representations (Irreducible Cartesian Tensors, ICTs) (zaverkin2024higher; schutt2021equivariant) for $O(3)$. Spherical tensors serve as the starting point for EGNN based on irreducible representations. However, they require Clebsch-Gordan transforms with significant computational overhead to derive desired irreducible representations from existing ones during each inference, making them computationally intensive. In contrast, ICTs are more efficient, as high-order representations can be easily computed using tensor products (specifically, outer products in EGNNs), making them more computationally friendly for most deep learning frameworks. The ICTs are obtained through the ICT decomposition matrices $H$,
${\rm vec}\left(T^{(n;j;q)}\right)=H^{(n;j;q)}{\rm vec}\left(T^{(n)}\right),$ (1)
where $T^{(n;j;q)}$ is an ICT, ${\rm vec}$ is the vectorization operation via flattening the tensor, $T^{(n)}$ is a given tensor, $n$ represents the rank of the tensor, and $j$ is called the weight (the order of irreducible representations, which is distinct from the learnable weights in neural networks) of the irreducible representation, and $q$ denotes the label (since the multiplicity for given $n$ and $j$ is sometimes greater than $1$). The decomposition matrices $H^{(n;j;q)}$ should be constructed such that $T^{(n;j;q)}$ are irreducible. A formal definition is introduced in Section LABEL:sec:ictdecomp. The challenge lies in the fact that $H$ of rank $n>5$ are unknown. The recent development of EGNNs research benefits from the irreducible decomposition matrices $H$ obtained from the physics and chemistry communities. The decomposition matrices for $n=4$ have been established by andrews1982irreducible. However, only very recently have we obtained the decomposition matrices for $n=5$ (bonvicini2024irreducible). Due to these limitations, state-of-the-art EGNNs cannot freely manipulate high-order ICTs or can only manipulate those whose weight equals their rank (zaverkin2024higher).
@@
Method Rank Orthogonality Time/Space complexity
@@
Table 1: Comparison with previous works. The complexity is discussed in terms of rank. The complexity of this work will be polynomial if we focus on the dimensions.
@@
This work has four main contributions:
  • We propose a method to efficiently calculate high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$, compared to predecessors, as shown in Table 1.
  • We propose a general parentage scheme to extend ICT decomposition into a generalized ICT decomposition applicable to arbitrary tensor product spaces.
@@
The following contents are organized as follows. In Section 1, we recall some basic ideas of representation theory, spherical harmonics, tensors, tensor product spaces, and we also introduce the Clebsch-Gordan coefficients and the parentage scheme. In Section LABEL:sec:problem, we formulate our problems: obtaining high-order ICT decomposition matrices and bases for arbitrary tensor product spaces. In Section LABEL:sec:highorderict, we introduce our theory and method to obtain high-order ICT decomposition matrices. In Section LABEL:sec:basis, we present the theory and method to obtain bases of equivariant spaces where the input and output spaces are the same. In Section LABEL:sec:extension, we extend the theory to the equivariant design spaces where the input and output spaces are different, and show the method to obtain bases for such general spaces. In Section LABEL:sec:relatedworks, we review related works on designing EGNNs and calculating ICT decomposition matrices. In Section LABEL:sec:discussion, we discuss the relation between pearce2023brauer and this work, as well as the theory for $SO(n)$, $O(n)$, and $SU(n)$. We provide concluding remarks in Section LABEL:sec:conclusion, and discuss the potential social impact in Section LABEL:sec:boardersocialimpact. We introduce the reproducibility details in Section LABEL:sec:repro. The acknowledgments are presented after the main text. The visualizations are shown in the appendix. A list of mathematical symbols is provided in Table LABEL:tab:symbols, and a mind map of this work is shown in Figure 1.
Table 2: Mathematical symbols and their meanings
Symbol Meaning Symbol Meaning
$\lVert\cdot\rVert_{col}$ The $L_{2}$-norm of an arbitrary column of $\cdot$ $G$ An arbitrary group
$SO(n)$ Special orthogonal group on $n$-dimensional space $SU(n)$ Special unitary group on $n$-dimensional space
$O(n)$ Orthogonal group on $n$-dimensional space $U(n)$ Unitary group on $n$-dimensional space
$\mathcal{S}$ A vector space $\rho_{\mathcal{S}}$ A group representation
$\mathcal{S}^{\otimes n}$ A rank $n$ tensor product space from $\mathcal{S}$ $N^{(n;j)}$ The multiplicity of the ICT of rank $n$ and weight $j$
$T$ An arbitrary tensor $G\to\mathcal{S}$ $T^{(n)}$ A Cartesian tensor of rank $n$
$\hat{T}^{(n;j;q;p)}$ An ICT of rank $n$, weight $j$, index $q$, and parity $p$ (sometimes omitted) $Q^{(l;p)}_{\mathcal{S}}$ The set of all possible paths to decompose a tensor in space $\mathcal{S}$
$H^{(n;j;q)}$ An ICT decomposition matrix to obtain ICTs of rank $n$, weight $j$, and index $q$ $\hat{H}^{(\mathcal{S};j;q)}$ A general ICT decomposition matrix of space $\mathcal{S}$, weight $j$ and index $q$
$C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}$ A CG coefficient with input weights $l_{1}$, $l_{2}$, output weight $l_{o}$, and the corresponding indices $m_{1}$, $m_{2}$, and $m_{o}$ (the angular-momentum quantum numbers in physics)
$C^{l_{1}l_{2}l_{o}}$ A CG tensor with input weights $l_{1}$, $l_{2}$, and output weight $l_{o}$ $\hat{C}^{l_{1}l_{2}l_{o}}$ A CG matrix by merging the first two dimensions of the original CG tensor
$P^{(path)}$ A path matrix generated by $path$, e.g., $(1\to 3\to 2)$ $\hat{P}^{(path)}$ A normalized path matrix generated by $path$
$D^{(\mathcal{S}\to\mathcal{S}^{\prime})}$ A change-of-basis matrix from space $\mathcal{S}$ to $\mathcal{S}^{\prime}$ $v^{(l;p)}$ A representation weight $l$ and parity $p$ irreducible representation acts on
${\rm Hom}_{G}$ The set of $G$-equivariant maps from one space to another ${\rm End}_{G}$ The set of $G$-equivariant maps from one space to itself
irr Returns all weights of irreducible representations in a decomposition ${\rm dim}$ The number of dimensions of a space
$(i_{1}\overset{k_{1}}{\to}i_{2}$ $\overset{k_{2}}{\to}\dots)$ A path where $k_{*}$ are bridge numbers (omitted if all being $1$) $(l,p)$ An irreducible representation space of weight $l$ and parity $p$
$\otimes$ Tensor product $\oplus$ Direct sum
$\odot$ Contraction ##\## Number of elements in a set
@@ 2 Preliminaries
2.1 Clebsch-Gordan coefficients
A review on irreducible representations and spherical spaces is in the appendix. The term weights in this paper refer to the order of irreducible representations. An important tool utilized in our work is the Clebsch-Gordan coefficients (CG coefficients). They are real numbers, and describe how to couple vectors from two spherical spaces into a new one. There are some rules for this coupling. First, given two vectors in weights $l_{1}$ and $l_{2}$ spherical spaces, the weights of the resulting vectors can only lie in $|l_{1}-l_{2}|\leq l_{o}\leq l_{1}+l_{2}$. For each index $m_{o}$ of the obtained vector $v_{l_{o}}\in(l=l_{o})$, it is calculated by homogeneous quadratic polynomials. Specifically, we have
$v^{(l_{o})}_{m_{o}}=\sum_{m_{o}=m_{1}+m_{2}}C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}v_{m_{1}}^{(l_{1})}v_{m_{2}}^{(l_{2})},$ (3)
and if we consider O(n) instead of SO(n), parity $p$ should also be considered. The parity is a property additional to the rank and weight, being $1$ or $-1$. Having $-1$ parity means that the representation will be inverse when the space is inverse, and vice versa. For example, the natural representations of O(n) always have $p=-1$. We will introduce the case that $p=1$ in the next section. Taking the parity into consideration, we can guarantee O(n) equivariance, and equation (3) becomes
$v^{(l_{o};p_{o})}_{m_{o}}=\mathbbm{1}_{(p_{o}=p_{1}p_{2})}\sum_{m_{o}=m_{1}+m_{2}}C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}v_{m_{1}}^{(l_{1};p_{1})}v_{m_{2}}^{(l_{2};p_{2})},$
where $\mathbbm{1}_{(exp)}$ is the indicator function, outputting $1$ if $exp$ holds true, or 00 otherwise. Our discussion will be the same no matter we consider $O(n)$ or $SO(n)$. But since Section LABEL:sec:extension, we need to take care of the parity which will play some roles in the further discussion. The CG coefficients are highly sparse, since $C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}=0$ if $m_{o}\neq m_{1}+m_{2}$. The coefficients are identical to Wigner-3j symbols up to multiplicative factors, so we will adopt Wigner-3j for the implementation. The evaluation of CG coefficients is nearly $\mathcal{O}(1)$ (if we do not consider the few calculations for the closed-form formula, which is cheap enough). To obtain the coefficients of $O(3)/SO(3)$, one often leverages the relation between $SU(2)$ and $SO(3)$: $SU(2)$ is a double cover of $SO(3)$, so the CG matrices of $SU(2)$ and $SO(3)$ are same up to a change-of-basis, from complex to real. The closed-form formulae for $SU(2)$ CG coefficients are already known. Thus, this process can be efficiently evaluated without obvious overhead.
Figure 1: The mind map of our method. The blocks in orange color are aligned with the three contributions of this work.
2.2 Tensors and tensor product spaces
To formally define a tensor, we need to consider vector space $V$ and covector space $V^{*}$. A tensor product is a bilinear map $\otimes:\mathcal{S}\times\mathcal{S}^{\prime}\to\mathcal{S}\otimes\mathcal{S}^{\prime}$. The basis of $\mathcal{S}\otimes\mathcal{S}^{\prime}$ is the set of $\hat{x}_{i}\otimes\hat{y}_{j}\in\mathcal{S}\otimes\mathcal{S}^{\prime}$, where $\hat{x}_{i}$ and $\hat{y}_{j}$ are basis vectors of $\mathcal{S}$ and $\mathcal{S}^{\prime}$, respectively. That is why one implements outer product as tensor product, as it meets the definition of tensor product. A tensor $T:V^{\otimes r}\otimes V^{*\otimes s}\to F$, where $F$ is some field, is a mathematical object that converts things between different tensor product spaces. For example, the representation $\rho(g)$ is a $(r=1,s=1)$-tensor that receives a vector $v$ which is a $(r=1,s=0)$-tensor, and output another $(r=1,s=0)$-tensor $\rho(g)v$, as $r$ and $s$ vanish correspondingly. Tensor product can help us build up the ranks of tensors, e.g., $V\otimes V\otimes V$ is a $(r=3,s=0)$ space. The representation space is accordingly $(r=3,s=3)$. Conversely, contraction can reduce the ranks of tensors. We express these operations via some index notations. For example, $T_{i_{1}i_{2}}=T^{\prime}_{i_{1}}T^{\prime\prime}_{i_{2}}$ is a tensor product, and $T_{i_{1}}=\sum_{j_{1}}T^{\prime}_{i_{1}j_{1}}T^{\prime\prime}_{j_{1}}$ is a contraction. Obviously, such operations can be mixed together, and we often omit the index of the output tensors and agree $i_{k}$ will be placed at the $k$th index of the output tensor by default. We also agree that $j_{1}$ and $j_{2}$ will vanish due to the contraction. We will not use Einstein summation convention in this work to keep the conventions of machine learning community, and to avoid ambiguous during derivations. One should note that contractions can be expressed in term of matrix representation. Consider $T=\sum_{j_{1},j_{2}}T^{\prime}_{i_{1}j_{1}j_{2}}T^{\prime\prime}_{j_{1}j_{2}}$, we have
$\begin{bmatrix}T_{1}\\ T_{2}\end{bmatrix}=\begin{bmatrix}T^{\prime}_{111}&T^{\prime}_{112}&T^{\prime}_{121}&T^{\prime}_{122}\\ T^{\prime}_{211}&T^{\prime}_{212}&T^{\prime}_{221}&T^{\prime}_{222}\end{bmatrix}\begin{bmatrix}T^{\prime\prime}_{11}\\ T^{\prime\prime}_{12}\\ T^{\prime\prime}_{21}\\ T^{\prime\prime}_{22}\end{bmatrix}.$
It is useful when we need to obtain the inverse of some linear operations. Also, contractions and tensor products are both $O(n)$-equivariant.
2.3 CG matrices and the parentage scheme
Figure 2: The parentage scheme. The numbers in the circle represent multiplicity. The sum of each column equals equation (LABEL:eq:multiplicity).
The $O(3)$ group has its representations in tensor product space $(\mathbbm{R}^{3})^{\otimes n}$, which are some $3^{n}\times 3^{n}$ matrices, or tensor product of $n$ $(r=1,s=1)$-tensors, which depends on one’s view. If we interpret them as matrices, then the objects they act on are simply $3^{n}$-dimensional vectors. The tensor product spaces $(\mathbbm{R}^{3})^{\otimes n}$ for $n>1$ are generally reducible. Recall CG coefficients, two irreducible spherical vectors $v$ and $v^{\prime}$ coupling can also be viewed as that they first form a rank 2 tensor, and then decompose into different new spherical vectors. These operations can also be viewed as a product of matrices and vectors. In this context, the CG coefficients $\hat{C}^{l_{1}l_{2}l_{o}}$ form a $(2l_{1}+1)(2l_{2}+1)\times(2l_{o}+1)$ matrix, which we call a CG matrix. If we reshape it to a $(2l_{1}+1)\times(2l_{2}+1)\times(2l_{o}+1)$ tensor, then we instead call it an $(l_{1},l_{2},l_{o})$-tensor $C^{l_{1}l_{2}l_{o}}$. Here are some important properties,
Lemma 1 (Orthogonality)
Given an $(l_{1},l_{2},l_{o})$-CG tensor $C^{l_{1}l_{2}l_{o}}$, we have
$\sum_{j_{1},j_{2}}C^{l_{1}l_{2}l_{o}}_{j_{1}j_{2}k}C^{l_{1}l_{2}l_{o}}_{j_{1}j_{2}k^{\prime}}=\sum_{j_{1},j_{2}}C^{l_{1}l_{2}l_{o}}_{j_{1}kj_{2}}C^{l_{1}l_{2}l_{o}}_{j_{1}k^{\prime}j_{2}}=\sum_{j_{1},j_{2}}C^{l_{1}l_{2}l_{o}}_{kj_{1}j_{2}}C^{l_{1}l_{2}l_{o}}_{k^{\prime}j_{1}j_{2}}=0$
for $k\neq k^{\prime}$. More generally, different CG tensors with the same $l_{1}$ and $l_{2}$, $l_{1}$ and $l_{3}$, or $l_{2}$ and $l_{3}$ are also orthogonal in this way.